# Solving Optimal Power Flow on a Data-Budget: Feature Selection on Smart Meter Data

### v1

@@ Abstract
How much data is needed to optimally schedule distributed energy resources (DERs)? Does the distribution system operator (DSO) have to precisely know load demands and solar injections at each bus of the feeder to solve an optimal power flow (OPF)? This work exploits redundancies in OPF’s structure and data to avoid communicating such data deluge, and explores the trade-off between data compression and grid’s performance. We propose an OPF data distillation framework involving two steps. The DSO first collects OPF data from only a subset of nodes. The DSO subsequently reconstructs the complete OPF data from the partial ones, and feeds them into the OPF solver. Selecting and reconstructing OPF data may be performed to either maximize the fidelity of reconstructed OPF data, or maximize the fidelity of OPF solutions corresponding to reconstructed data. Under the first objective, OPF data distillation is posed as a sparsity-regularized convex problem. Under the second objective, it is posed as a sparsity-regularized bilevel program. Both problems are solved using accelerated proximal gradient (PGD) algorithms. Numerical tests corroborate that the bilevel formulation enhances fidelity and feasibility of reconstructed OPF solutions, and that OPF solutions can be approximated reasonably well even from partial data.
@@ I Introduction
With the rampant integration of DERs, distribution system operators (DSOs) need to dispatch resources optimally at increasingly finer spatiotemporal scales. Solving the OPF requires collecting a large amount of data in near real-time. Currently, smart meters report packetized load demand data hourly or daily. Nonetheless, to control DERs effectively, the DSO may have to communicate with all customers and collect their active and reactive power demands every few minutes. This technical specification may be hard to meet due to communication, data privacy, and cyber-attack considerations. This work leverages redundancies in OPF features and the structure of the OPF to approximately solve the OPF using only a carefully selected subset of load demands.
@@
In transmission systems, the operator estimates voltages from grid measurements, infers loads by plugging voltages into the power flow equations, and feeds the inferred load values into real-time market operations. This workflow may not be replicated in real-time distribution systems operations due to a lack of observability at the grid edge. Under this setting, can DSOs control DERs based on partial grid data? The OPF can be viewed as a parametric optimization problem: Given load demands at all buses captured by vector $\boldsymbol{\theta}$, the OPF returns the optimal DER dispatch denoted by vector $\mathbf{x}(\boldsymbol{\theta})$. A key challenge in real-time DSO operations is that $\boldsymbol{\theta}$ may be uncertain or incomplete. The OPF mapping $\boldsymbol{\theta}\rightarrow\mathbf{x}$ has been extensively studied in the literature; see [1, 2] and references therein. Uncertainty quantification studies how a probabilistic characterization postulated on $\boldsymbol{\theta}$ propagates through the OPF [3]. Sensitivity analysis computes the Jacobian matrix of $\mathbf{x}$ with respect to $\boldsymbol{\theta}$; see [4, Ch. 6]. Stochastic or robust programming aims at securing system’s performance when $\boldsymbol{\theta}$ is not precisely known [5]. Reference [6] studies the value of data in grid operations and how obfuscation to preserve privacy can affect OPF solutions. Schemes for securing customer’s private load demand information are devised in [7].
@@
The focus here is on compressing $\boldsymbol{\theta}$ to approximately solve the OPF. Kron/Ward reduction techniques from the 1950’s and more recent ones aim at compressing the OPF problem itself via a reduced network model [8], [9], [10]. However, mapping results back to the actual system is non-trivial. Reference [11] applies principal component analysis (PCA) to a dataset of OPF solutions $\mathbf{x}$ so that a learning-to-optimize deep neural network (DNN) is trained to learn only a limited subspace of $\mathbf{x}$. Instead of $\mathbf{x}$, we aim to compress $\boldsymbol{\theta}$. References [12] and [13] suggest a joint learning and optimization approach, according to which OPF data are unknown and can only be predicted from primitive features $\boldsymbol{\phi}$, such as weather forecasts, historical trends, or partial OPF data. Rather than first inferring $\boldsymbol{\theta}$ from $\boldsymbol{\phi}$ and then mapping $\boldsymbol{\theta}$ to $\mathbf{x}$ via the OPF or an ML model, [12] and [13] advocate learning the mapping from $\boldsymbol{\phi}$ to $\mathbf{x}$. In our setting, the DSO has no access to $\boldsymbol{\phi}$ and needs to decide which OPF data should be sampled in real time. Keeping the OPF module in the workflow ensures that data distillation does not affect the OPF solver so that the same OPF solver is reused regardless of whether the DSO collects more or less data.
@@
The task of selecting relevant OPF data is reminiscent of sparse learning or feature selection in machine learning [14], wherein given feature vector $\boldsymbol{\theta}$, one tries to infer a target variable $\mathbf{x}(\boldsymbol{\theta})$ using only a subset of features. Although feature selection and reconstruction have been extensively studied in the ML context, we revisit them with a fresh look under the OPF lens. Clearly, selecting OPF data is closely related to the optimal placement of meters [15, 16], or actuators [17]. Rather than improving the grid’s observability or controllability, we select data to support OPF operations.
@@
    If the goal is to improve the fidelity of OPF data, we adopt a greedy algorithm based on the discrete empirical interpolation method (DEIM) and a proximal gradient algorithm for solving a group-lasso-type convex optimization problem (Section III); and
@@
    If the goal is to improve the fidelity of OPF solutions, we put forth a bilevel optimization program and develop a proximal gradient algorithm for solving it (Section IV).
The developed methods are evaluated on a single-phase simplification of the IEEE 37-bus benchmark feeder and the IEEE 123-bus multiphase feeder. We solved prototypical examples of the OPF that control reactive power injections from DERs to minimize losses while maintaining bus voltages within allowable limits. Grid loading conditions were simulated using a combination of synthetic and real-world load, solar, and electric vehicle data. The tests corroborate that fidelity improves when selecting more OPF features, OPF-aware data distillation yields better results in terms of approximating optimal DER schedules and complying with constraints, and OPF solutions can be approximated reasonably well given only a relatively small subset of OPF data. Conclusions and exciting open research directions are presented in Section VI. Partial preliminary results have been reported in the conference precursor of this work [18].
@@ II Problem Statement
A distribution system operator (DSO) wants to decide the optimal setpoints for reactive power injections by inverter-interfaced DERs. To this end, the operator routinely collects data across a feeder, which are subsequently used as parameters in an OPF. Such data may be of sizable volume as they include active injections by DERs as well as active and reactive demands from inflexible loads. Such data may be aggregated at the level of primary network buses. Under limited communication and/or sensing capabilities, the operator may not be able to possess all such data promptly to solve the OPF in real-time. In other words, if the OPF has to be solved every few minutes, the operator may be unable to communicate and read out data from every single bus of the primary network. A more realistic setup is that the operator can afford to install sensors only at $K$ out of the $N$ buses of a feeder. Hence, in real-time, the operator would communicate only with these $K$ sensors and rely solely on their data feeds to solve the OPF. Under this setup, two questions arise: i) How to sample data optimally? and ii) How to solve the OPF using partial data?
To address these questions, we deal with a prototypical OPF version. Consider the task of optimally deciding the reactive power setpoints by DERs. Given re/active loads and active solar generation on each bus, the DSO intends to decide optimally on DER reactive power setpoints. Such setpoints could be found by solving an OPF that minimizes ohmic power losses on distribution lines while ensuring bus voltages lie within allowable limits. The OPF considers a linearized grid model of a single-phase radial feeder. Although the technical exposition considers this stylized version of the OPF, our proposed methodology applies readily to other settings; Section V includes numerical tests on a multiphase feeder.
@@
Figure 1: (a) OPF mapping: The OPF as a parametric optimization problem. The OPF maps input data $\boldsymbol{\theta}\in\mathbb{R}^{P}$ to minimizers $\mathbf{x}\in\mathbb{R}^{N}$. (b) OPF data distillation: The operator first selects $K$ out of $P$ data to form $\tilde{\boldsymbol{\theta}}\in\mathbb{R}^{K}$. The entries of $\tilde{\boldsymbol{\theta}}$ are communicated from each selected bus to the operator over a communication network. The operator uses $\tilde{\boldsymbol{\theta}}$ to reconstruct an OPF parameter vector $\hat{\boldsymbol{\theta}}$ of length $P$, equal to the original parameter vector $\boldsymbol{\theta}$. The operator feeds $\hat{\boldsymbol{\theta}}$ into the OPF solver to find the DER dispatch $\mathbf{x}(\hat{\boldsymbol{\theta}})$. The OPF solver remains the same. How to optimally design the selection function (green) and the reconstruction function (blue)? These two functions are designed once and offline, and then used in real time. Should they be designed to minimize $\|\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}\|$ or $\|\mathbf{x}(\boldsymbol{\theta})-\mathbf{x}(\hat{\boldsymbol{\theta}})\|$? See Sections III and IV, respectively.
To solve (5), the DSO needs to know $\boldsymbol{\theta}$ in nearly real-time. This requirement may be impractical due to communication or privacy concerns. As a remedy, the operator may install sensors reading injections on a subset $K$ out of $P$ data sources. If a bus is instrumented with a sensor, it is reasonable to assume the operator has real-time access to the corresponding entries of $(\mathbf{p}_{g}-\mathbf{p}_{\ell})$ and $\mathbf{q}_{\ell}$. The pertinent questions are:
@@
This task of OPF data distillation is visualized in Fig. 1(b). The operator selects $K$ out of $P$ entries of $\boldsymbol{\theta}$ to create vector $\tilde{\boldsymbol{\theta}}$. Given the selected OPF data $\tilde{\boldsymbol{\theta}}$, the DSO reconstructs a data vector $\hat{\boldsymbol{\theta}}$ to feed into the OPF. The reconstructed vector $\hat{\boldsymbol{\theta}}$ should be of the original dimension $P$. This ensures backward compatibility with the OPF solver and keeps the solver independent of the data distillation scheme.
@@
Matrices $(\mathbf{C},\mathbf{S})$ can be selected to optimize a meaningful metric. One idea is to seek $(\mathbf{C},\mathbf{S})$ so $\hat{\boldsymbol{\theta}}$ is a good approximation of $\boldsymbol{\theta}$. Such distillation schemes focus on the fidelity of OPF data and will be henceforth termed Type-1 schemes. Taking an end-to-end vantage point, a second idea seeks $(\mathbf{C},\mathbf{S})$ so that $\mathbf{x}(\hat{\boldsymbol{\theta}})$ is a good approximant of $\mathbf{x}(\boldsymbol{\theta})$. Such schemes focus on the effect of distilled data on OPF decisions and will be termed Type-2 schemes. Type-1 schemes are developed in Section III and Type-2 schemes in Section IV.

### v2

@@ Abstract
How much data is needed to optimally schedule distributed energy resources (DERs)? Does the distribution system operator (DSO) have to know load demands at each bus of the feeder to solve an optimal power flow (OPF)? This work exploits redundancies in OPF’s structure and data to minimize the communication of such a data deluge, and explores the trade-off between data compression and the grid’s performance. We propose an OPF data distillation framework involving two steps: The DSO first collects OPF data from only a subset of nodes. It subsequently reconstructs the complete OPF data from the partial ones, and feeds them into the OPF solver. Selecting and reconstructing OPF data may be performed to maximize the fidelity of the reconstructed data or the associated OPF solutions. Under the first objective, OPF data distillation is posed as a sparsity-regularized convex problem. Under the second objective, it is posed as a sparsity-regularized bilevel program. Both problems are solved using proximal gradient algorithms. The second objective is superior in approximating OPF solutions at the expense of increased complexity. Numerical tests show that it enhances the fidelity and feasibility of the reconstructed OPF solutions, which can be approximated reasonably well even from partial data.
@@ I Introduction
With the rampant integration of DERs, distribution system operators (DSOs) need to dispatch resources optimally at increasingly finer spatiotemporal scales. Solving the OPF requires collecting a large amount of data in near real-time. Currently, smart meters report packetized load demand data hourly or daily. Nonetheless, to control DERs effectively, the DSO may have to communicate with all customers and collect their active and reactive power demands every few minutes. This technical specification may be challenging to meet due to concerns related to communication, data privacy, and cyberattacks. This work leverages redundancies in OPF features and the structure of the OPF to approximately solve the OPF using only a carefully selected subset of load demands. The primary objective is to reduce the real-time communication overhead between customers and the DSO. This is critical because although smart meters collect data at fine temporal scales, such data often arrives at the DSO’s control room in batches after a few hours due to networking issues.
@@
In transmission systems, the operator estimates voltages from grid measurements, infers loads by plugging voltages into the power flow equations, and subsequently feeds the inferred load values into real-time market operations. This workflow may not be replicated in real-time distribution system operations due to a lack of observability at the grid edge. Under this setting, can DSOs control DERs based on partial grid data? The OPF can be viewed as a parametric optimization problem: Given load demands at all buses captured by vector $\boldsymbol{\theta}$, the OPF returns the optimal DER dispatch denoted by vector $\mathbf{x}(\boldsymbol{\theta})$. A key challenge in real-time DSO operations is that $\boldsymbol{\theta}$ may be uncertain or incomplete. The OPF mapping $\boldsymbol{\theta}\rightarrow\mathbf{x}$ has been extensively studied in the literature; see [1, 2] and references therein. Uncertainty quantification studies how a probabilistic characterization postulated on $\boldsymbol{\theta}$ propagates through the OPF [3]. Sensitivity analysis computes the Jacobian matrix of $\mathbf{x}$ with respect to $\boldsymbol{\theta}$; see [4, Ch. 6]. Stochastic or robust programming aims at securing the system’s performance when $\boldsymbol{\theta}$ is not precisely known [5]. Reference [6] studies the value of data in grid operations and how obfuscation to preserve privacy can affect OPF solutions. Schemes for securing privacy in smart meter data are devised in [7, 8, 9]. The concept of differential privacy is adopted in [10] to solve the OPF while ensuring the privacy of electricity customers.
The focus here is on compressing $\boldsymbol{\theta}$ to approximately solve the OPF. Kron/Ward reduction techniques from the 1950’s and more recent ones aim at compressing the OPF problem itself via a reduced network model [11], [12], [13]. However, mapping results back to the actual system is non-trivial. Compression has also been leveraged in the context of learning-to-optimize, wherein a deep neural network (DNN) is trained to predict OPF solutions. References [14] and [15] suggest compressing the input and output space of the OPF mapping, respectively, so that the OPF mapping can be learned by a DNN having fewer trainable weights. Reference [16] introduces an encoder-decoder DNN structure to design a semi-centralized DER control policy. Although jointly trained, the encoder is implemented at the DSO and the decoder at DERs. Because the encoder output is of limited dimension, DERs can be controlled by a data-frugal control signal issued by the DSO.
@@
Instead of compressing $\boldsymbol{\theta}$ or $\mathbf{x}$, this work aims at selecting a subvector of $\boldsymbol{\theta}$ to feed into the standard OPF module. References [17] and [18] suggest a joint learning and optimization approach, according to which OPF data are unknown and can only be predicted from primitive features $\boldsymbol{\phi}$, such as weather forecasts, historical trends, or partial OPF data. Rather than first inferring $\boldsymbol{\theta}$ from $\boldsymbol{\phi}$ and then mapping $\boldsymbol{\theta}$ to $\mathbf{x}$ via the OPF or an ML model, references [17] and [18] advocate learning the mapping from $\boldsymbol{\phi}$ to $\mathbf{x}$. In our setting, the DSO has no access to $\boldsymbol{\phi}$ and must decide which OPF data to sample in real time. Keeping the OPF module in the workflow ensures that data distillation does not affect the OPF solver, allowing the same OPF solver to be reused regardless of whether the DSO collects more or less data.
@@
The task of selecting relevant OPF data is reminiscent of sparse learning or feature selection in machine learning [19], wherein given feature vector $\boldsymbol{\theta}$, one tries to infer a target variable $\mathbf{x}(\boldsymbol{\theta})$ using only a subset of features. Although feature selection and reconstruction have been extensively studied in the ML context, we revisit them with a fresh look under the OPF lens. Clearly, selecting OPF data is closely related to the optimal placement of meters or actuators. Reference [20] considers the task of inferring power injections at all buses of a distribution grid by polling smart meter data from only a subset of buses. The problem of installing actuators to improve controllability in transmission systems has been studied in [21]. Reference [22] activates a subset of inverter-based resources to learn the distribution grid topology, and [23] monitors a subset of buses in real time to ensure voltages remain within allowable limits. Perhaps most closely related to this work, reference [24] considers selecting a subset of smart meters to poll in real time to improve distribution grid state estimation. Rather than improving the grid’s observability or controllability, the goal here is to identify which data streams are most critical for approximating OPF solutions.
@@
    If the goal is to improve the fidelity of reconstructed OPF data, we adopt a greedy algorithm based on the discrete empirical interpolation method (DEIM) and a proximal gradient algorithm for solving a group-lasso-type convex optimization problem (Section III); and
@@
    If the goal is to improve the fidelity of reconstucted OPF solutions, we put forth a bilevel optimization program and develop a proximal gradient algorithm for solving it (Section IV).
The developed methods are evaluated on a single-phase simplification of the IEEE 37-bus benchmark feeder and a 1,136-bus feeder. We solved prototypical examples of the OPF that control reactive power injections from DERs to minimize losses while maintaining bus voltages within allowable limits. Grid loading conditions were simulated using a combination of synthetic and real-world load, solar, and electric vehicle data. The tests corroborate that fidelity improves when selecting more OPF features, OPF-aware data distillation yields better results in terms of approximating optimal DER schedules and complying with constraints, and OPF solutions can be approximated reasonably well given only a relatively small subset of OPF data. Conclusions and open research directions are presented in Section VI. Partial preliminary results have been reported in the conference precursor of this work [25].
Notation: Column vectors (matrices) are denoted by lower-(upper-) case letters. Symbol $(\cdot)^{\top}$ stands for transposition; $\mathbf{I}_{N}$ is the $N\times N$ identity matrix; $\|\mathbf{x}\|_{2}$ is the $\ell_{2}$-norm of vector $\mathbf{x}$; and $\|\mathbf{X}\|_{F}$ is the Frobenius matrix norm.
@@ II Problem Statement
A distribution system operator (DSO) wants to decide the optimal setpoints for reactive power injections by inverter-interfaced DERs. To this end, the operator routinely collects data across a feeder, which is subsequently used as parameters in an OPF. Such data may be of a sizable volume, as they include active injections by DERs, as well as active and reactive demands from inflexible loads. Such data may be aggregated at the level of primary network buses. Under limited communication and/or sensing capabilities, the operator may not be able to possess all such data promptly to solve the OPF in real-time. In other words, if the OPF must be solved every few minutes, the operator may be unable to communicate and read data from every single bus of the primary network. A more realistic setup is that the operator can afford to install sensors only at $K$ out of the $N$ buses of a feeder. Hence, in real-time, the operator would communicate only with these $K$ sensors and rely solely on their data feeds to solve the OPF. Under this setup, two questions arise: i) How to sample data optimally? and ii) How to solve the OPF using partial data?
To address these questions, we deal with a prototypical OPF version. Consider the task of optimally deciding the reactive power setpoints by DERs. Given re/active loads and active solar generation on each bus, the DSO intends to decide optimally on DER reactive power setpoints. Such setpoints could be determined by solving an OPF that minimizes ohmic power losses on distribution lines while ensuring that bus voltages remain within allowable limits. The OPF considers a linearized grid model of a single-phase radial test feeder. Although the technical exposition considers this stylized version of the OPF, our proposed methodology applies readily to other settings.
@@
Figure 1: (a) OPF mapping: The OPF as a parametric optimization problem. The OPF maps input data $\boldsymbol{\theta}\in\mathbb{R}^{P}$ to minimizers $\mathbf{x}\in\mathbb{R}^{N}$. (b) OPF data distillation: The operator first selects $K$ out of $P$ data to form $\tilde{\boldsymbol{\theta}}\in\mathbb{R}^{K}$. The entries of $\tilde{\boldsymbol{\theta}}$ are communicated from each selected bus to the operator over a communication network. The operator uses $\tilde{\boldsymbol{\theta}}$ to reconstruct an OPF parameter vector $\hat{\boldsymbol{\theta}}$ of length $P$, equal to the original parameter vector $\boldsymbol{\theta}$. The operator feeds $\hat{\boldsymbol{\theta}}$ into the OPF solver to find the DER dispatch $\mathbf{x}(\hat{\boldsymbol{\theta}})$. The OPF solver remains the same. How to optimally design the selection function (green) and the reconstruction function (blue)? These two functions are designed and implemented offline and then used in real-time. Should they be designed to minimize $\|\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}\|$ or $\|\mathbf{x}(\boldsymbol{\theta})-\mathbf{x}(\hat{\boldsymbol{\theta}})\|$? See Sections III and IV, respectively.
To solve (5), the DSO needs to know $\boldsymbol{\theta}$ in nearly real-time. This requirement may be impractical due to communication or privacy concerns. As a remedy, the operator may install sensors that read injections on a subset $K$ of $P$ data sources. If a bus is instrumented with a sensor, it is reasonable to assume the operator has real-time access to the corresponding entries of $(\mathbf{p}_{g}-\mathbf{p}_{\ell})$ and $\mathbf{q}_{\ell}$. The pertinent questions are:
@@
This task of OPF data distillation is visualized in Fig. 1(b). The operator selects $K$ out of $P$ entries of $\boldsymbol{\theta}$ to create vector $\tilde{\boldsymbol{\theta}}$. Given the selected OPF data $\tilde{\boldsymbol{\theta}}$, the DSO reconstructs a data vector $\hat{\boldsymbol{\theta}}$ to feed into the OPF. The reconstructed vector $\hat{\boldsymbol{\theta}}$ should be of the original dimension $P$. This ensures backward compatibility with the OPF solver and keeps the solver independent of the data distillation scheme, which may change with time.
@@
Matrices $(\mathbf{C},\mathbf{S})$ can be designed to optimize a meaningful metric. A first idea is to design $(\mathbf{C},\mathbf{S})$ so that $\hat{\boldsymbol{\theta}}$ approximates $\boldsymbol{\theta}$ well. Such distillation schemes focus on the fidelity of OPF data and will be henceforth referred to as Type-1 schemes. A second idea is to design $(\mathbf{C},\mathbf{S})$ so that $\mathbf{x}(\hat{\boldsymbol{\theta}})$ approximates $\mathbf{x}(\boldsymbol{\theta})$ well. Such schemes focus on the fidelity of OPF decisions and will be referred to as Type-2 schemes. Type-1 schemes are developed in Section III and Type-2 schemes in Section IV. Before presenting the proposed data distillation schemes, three remarks are in order.
Remark 1.
If the ultimate goal for the DSO is to find near-optimal OPF decisions using partial data, Type-2 schemes are preferable over Type-1 alternatives. Nonetheless, as it will become clear later, Type-2 schemes are computationally more demanding than Type-1 schemes. The performance advantage of Type-2 over Type-1 schemes is expected to diminish for larger data compression ratios $K/P$. Specifically, as the ratio $K/P$ tends to unity, the reconstructed OPF data $\hat{\boldsymbol{\theta}}$ approaches the actual data $\boldsymbol{\theta}$, and thus, the approximate decisions $\mathbf{x}(\hat{\boldsymbol{\theta}})$ are expected to lie close to $\mathbf{x}(\boldsymbol{\theta})$. In such cases, Type-1 schemes may perform reasonably well, forgoing the complexity of Type-2 schemes. On the other hand, for smaller $K/P$ ratios, the reconstructed data $\hat{\boldsymbol{\theta}}$ may be away from $\boldsymbol{\theta}$, in which case the increased complexity of Type-2 schemes pays off in recovering better OPF decisions.
Remark 2.
The primary motivation for OPF data distillation is to reduce the real-time communication overhead on the uplink between electricity customers and the DSO’s control room. Through the proposed schemes, the DSO communicates with only $K$ customers instead of $P$. Nonetheless, the proposed schemes could also be used to improve data privacy and cybersecurity issues as outlined next. Regarding data privacy, the DSO may ask customers if they consent to their real-time demand being used for scheduling purposes and select to sample only from those customers. Alternatively, the DSO may use data distillation schemes to identify data-critical customers and solicit their approval for data usage by providing discounts. Regarding cybersecurity, data distillation schemes may be beneficial in pinpointing the most critical data streams to protect by implementing higher cybersecurity protocols.
Remark 3.
Although the OPF data distillation schemes proposed in the following sections refer to the specific OPF model in (5), the suggested methodology is quite general. It can be readily applied to various OPF models, including AC OPF variants in single- or multi-phase grids. More specifically, Type 1 schemes are invariant to the OPF model. Even though Type 2 schemes depend on the OPF model, the only requirement is to be able to differentiate through the OPF mapping and compute $\nabla_{\boldsymbol{\theta}}\mathbf{x}$. In previous works [1, 2, 26], we have shown that this Jacobian matrix can be easily computed for various OPF formulations, including linear or quadratic programs (LP/QP) relying on linearized grid models, and second-order cone or semidefinite programs (SOCP/SDP) relying on nonlinear AC grid models. Regardless of the model, computing $\nabla_{\boldsymbol{\theta}}\mathbf{x}$ entails solving a system of linear equations, granted the optimal primal/dual OPF solutions have already been computed.